OpenStudy (anonymous):

1000 mm is equal to A. 1000 meters. B. 100 meters. C. 10 meters. D. 1 meter.

hero (hero):

Hint: milli = 1/1000

OpenStudy (anonymous):

so it would be A right?

hero (hero):

no

OpenStudy (anonymous):

ohhh D

hero (hero):

yes D is correct

OpenStudy (anonymous):

::) thanks

OpenStudy (anonymous):

The metric system is easy to use because it is based on units of ten. How many centimeters are there in 100 meters? A. 10 B. 100 C. 1000 D. 10,000

hero (hero):

x cm = 100m Solve for x

OpenStudy (anonymous):

10 right?

hero (hero):

Are you guessing or did you actually solve for x?

hero (hero):

centi = 1/100

hero (hero):

\[x \space\text{cm} = 100\text{m} \\ \frac{x}{100}\text{m} = 100\text{m} \] meters is a variable so it can be divided out: \[\frac{x}{100} = 100\] \[x = 100 \times 100\]

OpenStudy (anonymous):

10000 right?

hero (hero):

Yup